Show that the function (10×2 = 20)

a) u = x4 − 6x2 y2 + y is a solution of the two-dimensional Laplace equation.

b) u = sin w ct sin wx is a solution of the one-dimensional wave equation.

Grade:

1 Answers

Aman Bansal
12 years ago

Dear Swapnil,

The Laplace equation in two independent variables has the form

\frac{d^2\psi}{dx^2} + \frac{d^2\psi}{dy^2} \equiv \psi_{xx} + \psi_{yy} = 0.

[edit]Analytic functions

The real and imaginary parts of a complex analytic function both satisfy the Laplace equation. That is, if z = x + iy, and if

f(z) = u(x,y) + iv(x,y),\,

then the necessary condition that f(z) be analytic is that the Cauchy-Riemann equations be satisfied:

u_x = v_y, \quad v_x = -u_y.\,

where ux is the first partial derivative of u with respect to x.

It follows that

u_{yy} = (-v_x)_y = -(v_y)_x = -(u_x)_x.\,

Therefore u satisfies the Laplace equation. A similar calculation shows that v also satisfies the Laplace equation.

Conversely, given a harmonic function, it is the real part of an analytic function, f(z) (at least locally). If a trial form is

f(z) = \varphi(x,y) + i \psi(x,y),\,

then the Cauchy-Riemann equations will be satisfied if we set

\psi_x = -\varphi_y, \quad \psi_y = \varphi_x.\,

This relation does not determine ψ, but only its increments:

d \psi = -\varphi_y\, dx + \varphi_x\, dy.\,

The Laplace equation for φ implies that the integrability condition for ψ is satisfied:

\psi_{xy} = \psi_{yx},\,

and thus ψ may be defined by a line integral. The integrability condition and Stokes theorem implies that the value of the line integral connecting two points is independent of the path. The resulting pair of solutions of the Laplace equation are called conjugate harmonic functions. This construction is only valid locally, or provided that the path does not loop around a singularity. For example, if r and θ are polar coordinates and

\varphi = \log r, \,

then a corresponding analytic function is

f(z) = \log z = \log r + i\theta. \,

However, the angle θ is single-valued only in a region that does not enclose the origin.

Best Of luck

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert